A 25.0 mL sample of an electroplating solution is analyzed for its sulfuric acid concentration. It takes 46.8 mL of the 0.661 M NaOH from the example above to neutralızed the sample. Find the Molarity

of the acid.

The problem statement gives you the volume and concentration of NaOH, so you can determine the number of moles of NaOH. A balanced chemical equation gives you the information needed to find

moles of sulfuric acid.

Again, start by determining the quantity of H2SO4 in moles. Analyze the problem. Write the reaction equation so that you can make the moles of reactant to moles of product conversion.



1
Expert's answer
2021-03-04T06:37:18-0500

n(NaOH) = C*V = 0.661M*0.0468L = 0.0309mol

2NaOH + H2SO4 = Na2SO4 + 2H2O

According to the equation

n(H2SO4) = 0.0309/2 = 0.01545 mol

C(H2SO4) = n/V = 0.01545mol/0.025L = 0.618 M